# SciPostLayoutTree: A Dataset for Structural Analysis of Scientific Posters

### v1

@@ C.2 Additional Challenging Examples
C.3 Supplement to Analysis of Reading Order Prediction Improvements by Beam Search
@@
ResNet-50 DRGG-BS 79.2 (+5.0) 91.0 (+1.8) 93.9 (+4.0) 87.4 (+5.8) 78.3 (+5.9) 61.7 (+1.7) 71.4 (+6.0) 67.0 (+7.8)
DRGG-BEBS 85.6 (+4.1) 94.3 (+0.4) 96.7 (+2.0) 92.8 (+2.3) 85.0 (+4.0) 76.7 (+0.0) 85.0 (+2.5) 79.8 (+5.5)
ResNet-50 DRGG-BS 79.2 (+5.0) 91.0 (+1.8)⋆⋆ 93.9 (+4.0) 87.4 (+5.8) 78.3 (+5.9) 61.7 (+1.7) 71.4 (+6.0) 67.0 (+7.8)
DRGG-BEBS 85.6 (+4.1) 94.3 (+0.4) 96.7 (+2.0) 92.8 (+2.3) 85.0 (+4.0) 76.7 (+0.0) 85.0 (+2.5) 79.8 (+5.5)
ViT DRGG-BS 84.3 (+2.4) 93.2 (+1.2)⋆⋆ 95.1 (+1.7) 91.4 (+2.9) 83.4 (+1.9) 68.3 (+10.0)⋆⋆ 71.9 (+3.3)⋆⋆ 77.7 (+5.0)
DRGG-BEBS 87.4 (+2.8) 95.2 (+0.3) 96.9 (+1.2) 93.3 (+2.1) 86.0 (+2.6) 78.3 (+3.3) 85.3 (+1.9) 81.2 (+6.1)
Swin DRGG-BS 85.2 (+4.1) 93.8 (+1.2)⋆⋆ 95.2 (+1.5) 91.7 (+2.3) 84.4 (+3.2) 61.7 (+5.0) 74.9 (+4.4)⋆⋆ 76.4 (+4.0)
DRGG-BEBS 87.1 (+2.3) 95.0 (-0.4) 97.0 (+1.0) 92.8 (+0.9)⋆⋆ 86.5 (+2.3) 76.7 (-3.3) 86.6 (+1.4) 80.5 (+3.6)
DiT DRGG-BS 82.1 (+3.0) 91.4 (+0.7) 94.2 (+1.6) 90.2 (+3.0) 82.6 (+3.7) 51.7 (-1.7) 67.0 (+2.7) 74.7 (+4.9)
DRGG-BEBS 85.0 (+2.6) 94.5 (+0.6) 96.5 (+1.5) 92.9 (+2.6) 84.4 (+3.0) 75.0 (-1.7) 85.3 (+3.3)⋆⋆ 78.8 (+5.0)
InternImage DRGG-BS 84.9 (+2.7) 93.1 (+0.1) 96.1 (+1.5) 92.1 (+2.1) 84.9 (+2.1) 68.3 (+5.0) 80.1 (+4.9) 77.2 (+4.6)
DRGG-BEBS 87.1 (+2.3) 94.6 (-0.1) 96.6 (+0.8) 92.9 (+1.3) 86.9 (+2.8) 78.3 (+3.3) 86.4 (+2.2)⋆⋆ 79.7 (+3.1)
@@
ResNet-50 DRGG-BS 91.4 (+4.0) 76.3 (+5.6) 61.5 (+6.6) 70.1 (+11.9) 68.1 (+8.7)⋆⋆ 77.2 (+9.4)
DRGG-BEBS 95.0 (+1.9) 83.6 (+4.1) 76.1 (+6.6) 84.6 (+5.0)⋆⋆ 84.8 (+5.1)⋆⋆ 87.4 (+2.4)
ViT DRGG-BS 93.3 (+2.0) 82.1 (+2.3) 70.1 (+3.8)⋆⋆ 83.1 (+7.5) 75.4 (+2.2) 82.7 (+3.1)
DRGG-BEBS 95.5 (+1.4) 85.1 (+3.2) 77.4 (+3.4)⋆⋆ 83.6 (+7.0) 86.2 (+3.6) 86.6 (+4.7)⋆⋆
Swin DRGG-BS 93.5 (+1.9) 82.8 (+3.2) 72.6 (+4.5) 80.6 (+6.0)⋆⋆ 73.2 (+2.9) 85.0 (+6.3)⋆⋆
DRGG-BEBS 95.5 (+1.0) 85.0 (+2.3) 77.1 (+1.9) 84.1 (+2.0) 85.5 (+0.7) 87.4 (+4.7)⋆⋆
DiT DRGG-BS 91.9 (+1.9) 80.1 (+3.7) 70.7 (+5.8) 75.1 (+1.5) 72.5 (+2.9) 83.5 (+2.4)
DRGG-BEBS 94.6 (+1.5) 83.9 (+3.7) 75.9 (+4.9) 83.1 (+4.0)⋆⋆ 84.1 (+3.6) 89.0 (+0.8)
InternImage DRGG-BS 94.0 (+1.4) 83.4 (+3.7) 75.0 (+2.8) 84.6 (+5.0)⋆⋆ 75.4 (+1.4) 84.3 (+5.5)⋆⋆
DRGG-BEBS 95.0 (+1.0) 85.0 (+2.2) 77.4 (+3.2)⋆⋆ 89.1 (+3.0) 84.8 (+2.2) 85.0 (+3.1)
C.4 Supplement to Analysis of Parent-Child Prediction Improvements by BBox Embedding
@@
ResNet-50 DRGG-BE 89.0 (+18.1) 91.2 (+12.3) 95.6 (+11.9) 90.4 (+19.0) 74.0 (+15.0) 88.6 (+22.7) 85.4 (+17.8) 77.6 (+20.5)
DRGG-BEBS 93.6 (+13.6) 95.2 (+8.9) 97.8 (+9.4) 93.3 (+13.2) 80.3 (+18.9) 90.9 (+18.2)⋆⋆ 87.6 (+13.5) 79.0 (+8.3)⋆⋆
ViT DRGG-BE 90.9 (+5.0) 93.6 (+4.8) 96.4 (+6.0) 92.7 (+10.7) 80.3 (+10.2)⋆⋆ 81.8 (+11.4) 88.1 (+24.9) 79.0 (+12.2)
DRGG-BEBS 93.8 (+5.1) 96.3 (+3.6) 98.3 (+5.5) 95.6 (+9.2) 82.7 (+5.5) 84.1 (+6.8) 87.6 (+19.5) 82.9 (+10.2)
Swin DRGG-BE 91.5 (+8.8) 93.9 (+5.8) 96.7 (+7.3) 90.6 (+9.6) 76.4 (+6.3) 86.4 (+20.5) 89.7 (+24.9) 86.8 (+19.5)
DRGG-BEBS 94.1 (+7.3) 96.3 (+5.9) 98.2 (+6.6) 93.5 (+8.8) 77.2 (+7.9)⋆⋆ 86.4 (+13.6)⋆⋆ 89.7 (+21.1) 87.8 (+13.2)
DiT DRGG-BE 88.9 (+7.6) 92.8 (+7.0) 95.9 (+7.4) 91.7 (+10.9) 79.5 (+9.4)⋆⋆ 84.1 (+25.0) 84.9 (+32.4) 73.7 (+8.8)⋆⋆
DRGG-BEBS 92.6 (+7.8) 95.9 (+6.7) 97.6 (+7.2) 94.8 (+11.1) 85.8 (+15.7) 81.8 (+25.0) 87.0 (+30.3) 80.5 (+9.3)
InternImage DRGG-BE 89.6 (+5.1) 93.6 (+5.6) 96.9 (+6.7) 91.7 (+9.6) 77.2 (+3.9) 81.8 (+13.6) 84.9 (+15.1) 83.9 (+17.6)
DRGG-BEBS 93.5 (+6.0) 96.2 (+5.6) 98.2 (+6.1) 95.4 (+10.7) 81.9 (+6.3) 86.4 (+13.6)⋆⋆ 88.6 (+13.5) 88.8 (+16.1)
@@
ResNet-50 DRGG-BE 96.5 (+7.7) 87.8 (+17.1) 85.9 (+22.5) 87.4 (+21.7) 85.7 (+23.3) 88.3 (+32.7)
DRGG-BEBS 97.9 (+6.4) 92.4 (+12.9) 91.9 (+15.5) 91.1 (+13.7) 90.4 (+17.1) 93.7 (+28.3)
ViT DRGG-BE 97.3 (+5.5) 89.6 (+4.4) 89.7 (+8.7) 90.2 (+8.2) 86.0 (+8.7) 90.7 (+23.9)
DRGG-BEBS 98.4 (+4.9) 92.4 (+3.9) 93.8 (+6.7) 93.2 (+5.5) 92.9 (+9.6) 95.1 (+23.4)
Swin DRGG-BE 97.3 (+5.5) 90.6 (+8.6) 90.6 (+10.4) 88.9 (+9.7) 89.8 (+16.1) 91.2 (+26.8)
DRGG-BEBS 98.6 (+5.8) 93.1 (+7.2) 93.4 (+8.6) 91.5 (+7.7) 92.2 (+12.7) 94.1 (+20.5)
DiT DRGG-BE 96.6 (+7.5) 87.7 (+6.1) 88.1 (+10.8) 89.2 (+9.2) 87.9 (+15.2) 87.8 (+15.6)
DRGG-BEBS 97.9 (+7.2) 92.6 (+7.8) 91.6 (+9.1) 91.8 (+7.9) 91.9 (+15.2) 94.6 (+20.0)
InternImage DRGG-BE 97.0 (+4.8) 90.1 (+7.0) 89.1 (+8.3) 89.9 (+8.9) 88.5 (+12.7) 91.2 (+23.9)
DRGG-BEBS 98.5 (+5.2) 92.9 (+6.8) 93.0 (+8.2) 92.8 (+7.7) 94.1 (+13.4) 95.1 (+17.6)
C.5 Effect of Beam Width

### v2

@@ C.3 Additional Challenging Examples
C.4 Supplement to Analysis of Reading Order Prediction Improvements by Beam Search
@@
ResNet-50 DRGG-BS 79.4 (+5.0) 91.5 (+1.7)⋆⋆ 93.9 (+4.0) 87.7 (+6.1) 77.8 (+6.5) 58.3 (+0.0) 73.6 (+5.8) 67.4 (+8.2)
DRGG-BEBS 86.0 (+4.0) 94.5 (+0.6) 96.7 (+2.0) 92.7 (+2.4) 84.2 (+3.6) 73.3 (-1.7) 87.5 (+2.9) 79.8 (+5.5)
ViT DRGG-BS 84.6 (+2.9) 93.6 (+1.2)⋆⋆ 95.3 (+1.8) 91.3 (+2.6) 82.8 (+2.1) 63.3 (+5.0) 74.5 (+4.1) 78.0 (+5.2)
DRGG-BEBS 87.7 (+2.8) 95.2 (+0.0) 97.2 (+1.3) 93.1 (+2.0) 85.4 (+2.6) 75.0 (+5.0) 89.3 (+2.3) 81.4 (+5.7)
Swin DRGG-BS 85.6 (+4.0) 93.9 (+1.3) 95.2 (+1.4) 91.7 (+2.6) 83.5 (+3.2) 60.0 (+5.0) 77.1 (+4.1)⋆⋆ 77.2 (+4.7)
DRGG-BEBS 87.9 (+2.4) 95.0 (-0.5) 97.0 (+1.1) 92.7 (+1.0)⋆⋆ 85.5 (+2.8) 73.3 (-5.0) 89.0 (+2.0) 81.1 (+3.7)
DiT DRGG-BS 82.4 (+3.1) 91.7 (+0.8) 94.5 (+1.8) 90.1 (+3.4) 81.6 (+3.8) 50.0 (-5.0) 69.6 (+2.9) 75.8 (+5.9)
DRGG-BEBS 85.3 (+2.8) 94.6 (+0.5) 96.6 (+1.6) 92.9 (+2.7) 83.3 (+3.3) 71.7 (-3.3) 88.1 (+2.3) 79.3 (+5.5)
InternImage DRGG-BS 85.1 (+2.6) 93.4 (+0.2) 95.8 (+1.5) 92.1 (+2.4) 83.4 (+2.1) 63.3 (+3.3) 80.9 (+5.2) 77.8 (+4.7)
DRGG-BEBS 87.5 (+2.1) 94.8 (+0.0) 96.4 (+0.7) 92.9 (+1.3) 85.9 (+2.9) 73.3 (+3.3) 88.1 (+2.3)⋆⋆ 80.1 (+3.0)
@@
ResNet-50 DRGG-BS 91.3 (+4.1) 76.5 (+5.9) 62.1 (+6.7) 71.7 (+13.1) 71.0 (+9.9) 78.9 (+8.1)⋆⋆
DRGG-BEBS 94.7 (+1.9) 84.0 (+4.1) 77.4 (+6.4) 85.4 (+5.6)⋆⋆ 86.3 (+5.3)⋆⋆ 89.4 (+3.3)
ViT DRGG-BS 93.2 (+2.0) 82.6 (+2.7) 72.1 (+4.4) 83.8 (+8.6) 77.1 (+3.1) 85.4 (+3.3)
DRGG-BEBS 95.3 (+1.3) 85.5 (+3.1) 80.5 (+4.4) 86.4 (+8.1) 87.0 (+2.3) 89.4 (+4.9)⋆⋆
Swin DRGG-BS 93.3 (+2.0) 83.2 (+3.1) 74.1 (+4.7) 81.3 (+6.1)⋆⋆ 76.3 (+3.8) 87.8 (+4.9)
DRGG-BEBS 95.2 (+1.0) 85.5 (+2.6) 78.7 (+2.2) 84.3 (+3.0) 87.0 (+1.5) 90.2 (+4.9)⋆⋆
DiT DRGG-BS 91.8 (+2.0) 80.6 (+4.1) 72.5 (+6.0) 77.3 (+4.0) 74.8 (+3.1) 86.2 (+2.4)
DRGG-BEBS 94.4 (+1.6) 84.1 (+3.8) 77.2 (+5.1) 83.3 (+3.5) 87.8 (+3.8) 91.1 (+0.8)
InternImage DRGG-BS 93.7 (+1.5) 83.3 (+3.6) 74.9 (+3.1) 84.8 (+6.1) 76.3 (+1.5) 87.0 (+5.7)⋆⋆
DRGG-BEBS 94.8 (+0.9) 85.1 (+2.1) 78.0 (+4.0) 88.9 (+2.5) 86.3 (+2.3) 87.8 (+3.3)
C.5 Supplement to Analysis of Parent-Child Prediction Improvements by BBox Embedding
@@
ResNet-50 DRGG-BE 89.2 (+18.2) 91.4 (+12.1) 95.7 (+11.5) 90.2 (+19.5) 69.5 (+16.3) 88.4 (+18.6)⋆⋆ 86.5 (+18.0) 82.4 (+22.8)
DRGG-BEBS 93.3 (+13.3) 95.3 (+8.4) 98.0 (+9.2) 92.6 (+13.6) 74.5 (+19.1) 90.7 (+18.6)⋆⋆ 88.8 (+14.0) 83.4 (+10.4)⋆⋆
ViT DRGG-BE 91.3 (+5.2) 94.2 (+4.8) 96.6 (+5.9) 91.7 (+11.0) 75.2 (+12.1) 83.7 (+9.3) 89.9 (+26.4) 83.4 (+15.0)
DRGG-BEBS 94.3 (+5.3) 97.1 (+3.7) 98.4 (+5.5) 94.7 (+9.3) 77.3 (+7.1) 86.0 (+7.0) 89.9 (+20.8) 87.6 (+12.4)
Swin DRGG-BE 92.0 (+9.3) 94.6 (+6.1) 96.9 (+7.2) 89.4 (+9.1) 70.2 (+6.4)⋆⋆ 88.4 (+20.9) 91.6 (+24.7) 90.7 (+20.7)
DRGG-BEBS 94.8 (+7.9) 97.1 (+6.1) 98.5 (+6.6) 92.1 (+8.5) 73.0 (+9.9) 88.4 (+14.0)⋆⋆ 91.0 (+21.3) 90.2 (+14.0)
DiT DRGG-BE 89.2 (+7.6) 93.2 (+7.0) 96.0 (+7.4) 91.1 (+11.3) 73.0 (+9.9)⋆⋆ 88.4 (+25.6) 88.2 (+34.8) 78.2 (+9.8)
DRGG-BEBS 92.8 (+7.7) 96.3 (+6.5) 97.8 (+7.2) 94.7 (+12.1) 80.1 (+14.9) 83.7 (+23.3)⋆⋆ 90.4 (+32.6) 84.5 (+9.3)
InternImage DRGG-BE 89.9 (+5.2) 94.2 (+5.5) 97.2 (+6.7) 91.1 (+10.2) 70.9 (+4.3) 83.7 (+11.6) 87.6 (+16.3) 89.1 (+19.2)
DRGG-BEBS 93.8 (+6.1) 96.9 (+5.5) 98.5 (+6.1) 94.3 (+10.8) 75.9 (+7.1)⋆⋆ 88.4 (+11.6) 90.4 (+13.5) 92.2 (+16.6)
@@
ResNet-50 DRGG-BE 96.3 (+7.7) 88.0 (+17.4) 86.2 (+22.1) 88.7 (+21.4) 86.4 (+22.5) 88.3 (+32.2)
DRGG-BEBS 97.8 (+6.4) 92.2 (+13.1) 92.0 (+14.7) 92.3 (+13.3) 91.1 (+16.5) 93.7 (+28.3)
ViT DRGG-BE 97.2 (+5.7) 89.8 (+4.7) 90.7 (+8.7) 91.8 (+8.5) 87.0 (+8.2) 91.2 (+23.9)
DRGG-BEBS 98.3 (+5.1) 92.8 (+4.2) 94.8 (+6.6) 94.9 (+5.7) 94.0 (+9.5) 95.6 (+23.4)
Swin DRGG-BE 97.2 (+5.6) 90.5 (+8.5) 91.4 (+10.7) 90.7 (+9.7) 90.8 (+15.8) 91.7 (+26.8)
DRGG-BEBS 98.5 (+5.8) 93.4 (+7.6) 94.2 (+9.0) 93.5 (+7.9) 93.4 (+12.3) 94.6 (+20.5)
DiT DRGG-BE 96.5 (+7.5) 88.1 (+6.6) 88.7 (+10.8) 90.5 (+8.8) 88.6 (+14.9) 87.8 (+15.6)
DRGG-BEBS 97.9 (+7.2) 92.9 (+7.6) 92.4 (+9.5) 93.1 (+7.7) 93.0 (+15.2) 94.6 (+20.0)
InternImage DRGG-BE 97.0 (+4.8) 90.3 (+7.1) 90.0 (+8.9) 91.5 (+8.7) 89.9 (+12.7) 91.7 (+23.9)
DRGG-BEBS 98.4 (+5.2) 93.1 (+7.1) 93.5 (+8.0) 94.6 (+7.5) 95.6 (+13.3) 95.6 (+17.6)
C.6 Effect of Beam Width